OpenStudy (anonymous):

derivative: y=3^(xlnx)

hartnn (hartnn):

know chain rule ?

OpenStudy (anonymous):

somewhat

OpenStudy (anonymous):

would it be 3^(xlnx) times d/dx (xlnx)

hartnn (hartnn):

what is d/dx(a^x) ??

hartnn (hartnn):

\(d/dx (a^{f(x)}) = (a^{f(x)})ln \:a\quad d/dx(f(x))\)

hartnn (hartnn):

u just missed ln 3

OpenStudy (anonymous):

so ln3^(xlnx) times d/dx (xlnx) ?

hartnn (hartnn):

3^(xlnx) ln 3 times d/dx (xlnx)

OpenStudy (anonymous):

is that some type of shortcut?

hartnn (hartnn):

no, that is chain rule.

hartnn (hartnn):

since d/dx of a^x = a^x ln a from chain rule, we have \(d/dx (a^{f(x)}) = (a^{f(x)})ln \:a\quad d/dx(f(x))\)

hartnn (hartnn):

now just put here a=3, f(x) = x ln x

OpenStudy (anonymous):

so then we just need to use the product rule for the right side correct

hartnn (hartnn):

froduct rule for d/dx (x ln x) yes, correct.

OpenStudy (anonymous):

alryte got it thanks a lot hart

hartnn (hartnn):

welcome ^_^